OpenStudy (ineedhelpnow):

we plan to mix a-grade coffee that costs $9.50 per pound with b-grade coffee that costs $7.00 per pound to create a 20-pound blend that will sell for $8.50 per pound. how many pounds of each grade coffee are required?

OpenStudy (anonymous):

12 Pounds of A-Grade 8 Pounds of B-Grade

OpenStudy (ineedhelpnow):

how did you get that?

OpenStudy (anonymous):

Let us assume in that mixture of 20 pounds we will mix x pound of A grade. So, (20-x) will be of B-grade $ 9.5x + $ 7*(20-x) = $ 8.5*20 solve for x, 9.5x + 140 - 7x = 170 2.5x = 170-140 = 30 x = 30/2.5 = 12 Pound A-Grade 20-x = 8 pound B-Grade